# Local Surface Parameterizations via Geodesic Splines

### v1

@@ 4. Evaluation and Results
Table 1. Hyperparameter listings of all the results in the paper, where $m$ is the number of radial curves, $n$ is the number of steps, and $h$ is the step size.
Surface Fig. Type $m$ $n$ $h$ Smoothing Enabled?
Saddle 9 Analytic Implicit 50 10 $1.0\times 10^{-2}$ Yes
@@
Cone 12 SDF 50 12 $1.0\times 10^{-2}$ No
Bunny (coarse) 10 (top right) Mesh 20 20 $2.5\times 10^{-2}$ Yes
Bunny (fine) 10 (bottom right) Mesh 20 20 $2.5\times 10^{-2}$ Yes
@@
Brain 1 Mesh 50 10, 15, 20 $5.0\times 10^{-3}$ Yes
@@
Terrain 1 Point Cloud 50 10 $1.0\times 10^{-2}$ Yes
Star Torus 13 SDF 50 40 $1.0\times 10^{-2}$ Yes
Shellfish 11 (left) Neural Implicit 50 15 $1.0\times 10^{-2}$ Yes
Einstein 11 (middle) Neural Implicit 50 20 $5.0\times 10^{-3}$ Yes
Metratron 11 (right) Neural Implicit 50 15 $1.0\times 10^{-2}$ Yes
@@
Spot 14 Mesh 50 20 $1.0\times 10^{-2}$ Yes
@@ 4.1. Implementation
We implemented our code primarily in Python using the PyTorch library (Paszke et al., 2019) to have a framework for automatic differentiation, which we used to compute gradients of implicit functions and interface with trained neural implicits. We used a simple feature grid architecture for the neural implicits used in the paper, based on the NGLOD architecture (Takikawa et al., 2021) with a single $128^{3}$ feature grid, and an MLP with two 128-unit hidden layers and softplus activations. Point cloud implicits were obtained using a LogSumExp smooth distance function (Madan and Levin, 2022). Also, we used JIT compilation for analytic implicits and signed distance functions to speed up their evaluation. To have comparable parameters across all experiments, we isotropically rescaled every object to fit within the cube $[-1,1]^{3}$ (Table 1).
@@ 4.2. Performance
Asymptotically, each component of a single tracing step (projection, substepping, tridiagonal linear system solving) takes $O(m)$ time to advance the integration front from $\psi_{j}$ to $\psi_{j+1}$, so tracing $n$ steps is overall $O(mn$). Fitting the surface also takes $O(mn)$ time to build all of the $n$ isoline splines, and $O(n)$ time to evaluate the surface spline at a point (mainly to build $\tilde{\gamma}_{\theta}$). One key advantage of our approach is that the performance is output-sensitive; in other words, it does not strongly depend on the complexity or size of the underlying surface (though in some cases like triangle mesh distance, each implicit distance query is $O(\log N)$ where $N$ is the number of mesh triangles). For example, mesh-based heat methods (Sharp et al., 2019; Herholz and Alexa, 2019) are slightly superlinear in $N$ from solving sparse matrices, but if $mn\ll N$, this is much worse than $O(mn)$. In Table 2, we show a timing breakdown of the examples in Fig. 1, which are representative of the types of implicit functions we used throughout the paper. All timings were recorded on a 2020 MacBook Pro with an M1 processor. From these results, we see that none of the examples took more than 1.5s to complete, and the primary bottleneck is tracing, particularly root finding in each substep (line 8 of Alg. 1). Meshing the log map takes hundreds of milliseconds for every example, though this is because we evaluate the surface at over 10000 points to build the mesh, so this can easily be reduced if a coarser mesh is acceptable. Furthermore, it is apparent that overall performance is proportional to the amount of time required to evaluate the implicit $f$ and its gradient. Overall, these results are sufficient for interactive applications, even on the LiDAR terrain point cloud with over 29 million points, and suggest that speeding up the computation of $f$ and a faster root finding method can further improve performance.
@@ 4.3. Parameter Analysis
The hyperparameters $m$ (number of traced radial curves), $n$ (number of steps taken), and $h$ (step size) are the main quantities used to affect the accuracy of the parameterizations. We ran our method on a unit sphere (with no substepping and no holonomy smoothing) using a range of values for each of these hyperparameters (where $nh$, the total radius of the map in tangent space, is always $1$) and computed the average $L^{2}$ error of our output from the analytic exponential map, at 2000 samples inside the parameterized disc in tangent space and 500 points along the boundary of the disc (Fig. 7). Overall, we observe that both adding more steps (and correspondingly smaller substeps) and more radial curves can improve the results, but after enough steps/radial curves, adding more steps/radial curves does not seem to improve the results. In particular, increasing $m$ from 5 to 50 significantly reduces the error, though there is little difference between using 50 or 500 radial curves; as such, we use $m=50$ for most of the results in the paper. Furthermore, we see that the error levels out under refinement of $h$, but at different points when using 5 radial curves (10 steps) and 50/500 radial curves (100 steps). To further investigate the source of the error, we also plotted the error from our traced points along a single radial curve for each step size (since this is on a sphere, there is little difference between each individual radial curve). This error curve follows a very similar trend to the larger $m$ error curves which decrease and eventually level out around $10^{-5}$, which suggests that once there are sufficiently many traced radial curves, the primary source of error is the tracing rather than the spline interpolation. Since the error in the traced points accumulates over more steps, there is a tradeoff between accumulating larger per-step error over fewer steps, and accumulating smaller per-step error over more steps, which causes the error to level out under refinement. Despite the lack of error reduction beyond $10^{-5}$, this is quite small, and as demonstrated in the examples throughout the paper, does not cause practical issues.
@@
Holonomy smoothing strongly influences the output, and is controlled by a regularization parameter $\kappa$. In Fig. 8, we examine the effect of $\kappa$; as it decreases, the effect of the smoothing is diminished, and in the limit as $\kappa\to 0$, the smoothing is completely skipped. We built a local parameterization for a neural implicit representing the Stanford armadillo, around a point near a small indent in the forearm created by isosurface noise. With large $\kappa$ (i.e., strong smoothing), we see that the map has low distortion away from the origin, and although there is some distortion near the pothole itself, this distortion largely disappears as the map expands. Meanwhile, as $\kappa$ decreases, the map’s quality degrades; some regions fold over each other (seen through colormap discontinuities), while other regions significantly stretch away from the origin. As a result, we exclusively use $\kappa=10^{3}$ everywhere else in the paper.
@@ 4.4. Comparisons
Compared to prior local parameterization methods, our method can extract significantly more information out of implicit surfaces. Rather than simply sampling the surface for points, and using nearest neighbor information to, e.g., build graphs or Laplacians, we generate the samples we need through tracing, while also being able to evaluate the generated map on other surface points. To demonstrate this advantage, we generated a local parameterization around a saddle point with 500 traced samples, and rejection sampled the same number of surface points from a 3-dimensional ball around the map origin in $\mathbb{R}^{3}$ such that all the points are contained in our map. Then, we evaluated the log map of our method on those sample points, and also provided those points as input to both versions of discrete exponential maps (DEM) (Schmidt et al., 2006; Schmidt, 2013) and the vector heat method (Sharp et al., 2019). We used our own implementation for the DEM methods and the authors’ released code for the vector heat method. For each map, we built a mesh connecting the samples by using the connectivity from the Delaunay triangulation computed in tangent space. As seen in Fig. 9, all other methods produce high distortion away from the origin, and the vector heat method struggles throughout the entire map. It is worth noting that these struggles are exacerbated by the naïve point sampling method, which tends to produce clusters of points that can be prevented by more sophisticated blue noise sampling techniques (Witkin and Heckbert, 1994; De Goes et al., 2012); nevertheless, our method produces low distortion throughout the map despite the sample quality, since we do not depend on them for generating the parameterization.
@@
Figure 10. On a coarse bunny mesh (top), a local parameterization generated by our method (right) avoids the severe artifacts near the boundary of the map (left, centre left) and angular distortion near the origin (centre, centre right) seen by other methods. All methods improve after one level of Loop subdivision (bottom), but the DEM-based methods still possess boundary artifacts (left, centre left).
@@
Since we trace points along the surface, another interesting property of our method on triangle meshes in particular is that our result quality is independent of the underlying triangulation. In Fig. 10 we compare our method to both DEM methods, the vector heat method, and another heat-based method (Herholz and Alexa, 2019), for which we used our own implementation, taking care to use intrinsic Delaunay operators from libigl (Jacobson et al., 2018). For this experiment, we computed the 10-ring of an origin vertex (i.e., all vertices at most 10 edges away from the origin) on a coarse Stanford bunny mesh with 3500 vertices, and used these as the evaluation points for all methods. The DEM methods only used these points and their derived connectivity from the mesh, while we used the entire mesh to define the required operators for the heat methods (since we would change the boundary conditions if we only used the local patch). Our method does not build a map from pre-selected discrete points (aside from the map origin), so to make a best-effort comparison we ran our method so that the resulting map both contained the same 10-ring and used a comparable point budget (400 traced samples vs. 453 vertices in the 10-ring — see Table 1 for the parameters we used). The original mesh connectivity was used to connect the 10-ring evaluation points into a continuous patch for all methods. We see that our method produces by far the least distortion in the map, and just like in Fig. 9, the DEM methods exhibit distortion far away from the origin, while the heat-based methods struggle throughout the map, since the triangulation is not of a high enough quality to produce well-conditioned operators. After one level of Loop subdivision (which produces a mesh with roughly 14000 vertices), we ran the same experiment on the 20-ring to cover the same region on the surface (now containing 1720 vertices), and did not change the parameters in our method (though we should clarify that the underlying surface is now the subdivided mesh and not the original mesh). Now we see that our method produces virtually the same result as before, the heat-based methods perform much better with very little distortion, and the DEM methods improve but still exhibit artifacts at the boundary of the map. However, these other methods only achieve this quality with 4 times the number of points that we need to achieve similar quality, and cannot achieve this at all on a coarser mesh.
@@ 4.5.1. Local Texturing
A variety of examples of texturing different geometry types is provided in Fig. 1. We show several decals on each surface, where each decal is represented by a meshed version of the spline surface, and can see that none of the resulting images are distorted, while still conforming to the original surface. The terrain point cloud, owl neural implicit, and brain mesh all have challenging geometry with surface bumps, ridges, and divets that make parameterization more difficult. By placing maps in small regions, however, even regions like the owl’s plumage and the heavily folded surface of the brain can be textured.
@@
Figure 11. We obtain high-quality maps on a variety of neural implicit surfaces, even in the presence of noisy reconstruction artifacts and fine ground truth details.

### v2

@@ 4. Evaluation and Results
Table 1. Parameter listings of all the results in the paper, where $m$ is the number of radial curves, $n$ is the number of steps, and $h$ is the step size. The “Size” column lists the number of points in a point cloud, or the number of vertices in a mesh.
Surface Fig. Type Size $m$ $n$ $h$ Smoothing Enabled?
Saddle 10 Analytic Implicit 50 10 $1.0\times 10^{-2}$ Yes
@@
Cone 14 SDF 50 12 $1.0\times 10^{-2}$ No
Bunny (coarse) 12 (top right) Mesh 3,485 20 20 $2.5\times 10^{-2}$ Yes
Bunny (fine) 12 (bottom right) Mesh 13,934 20 20 $2.5\times 10^{-2}$ Yes
@@
Brain 1 Mesh 73,820 50 10, 15, 20 $5.0\times 10^{-3}$ Yes
@@
Terrain 1 Point Cloud 29,474,037 50 10 $1.0\times 10^{-2}$ Yes
Star Torus 15 SDF 50 40 $1.0\times 10^{-2}$ Yes
Shellfish 13 (left) Neural Implicit 50 15 $1.0\times 10^{-2}$ Yes
Einstein 13 (middle) Neural Implicit 50 20 $5.0\times 10^{-3}$ Yes
Metratron 13 (right) Neural Implicit 50 15 $1.0\times 10^{-2}$ Yes
@@
Spot 9 Mesh 2,930 50 20 $1.0\times 10^{-2}$ Yes
Shark 16 Mesh 10,054 50 20 $1.0\times 10^{-2}$ Yes
@@ 4.1. Implementation
We implemented our code primarily in Python using the PyTorch library (Paszke et al., 2019) to have a framework for automatic differentiation, which we used to compute gradients of implicit functions and interface with trained neural implicits. We used a simple feature grid architecture for the neural implicits used in the paper, based on the NGLOD architecture (Takikawa et al., 2021) with a single $128^{3}$ feature grid, and an MLP with two 128-unit hidden layers and softplus activations. Point cloud implicits were obtained using a LogSumExp smooth distance function (Madan and Levin, 2022). Also, we used JIT compilation for analytic implicits and signed distance functions to speed up their evaluation. To have comparable parameters across all experiments, we isotropically rescaled every object to fit within the cube $1,1]^{3}$ (Table 1).
@@ 4.2. Performance
Asymptotically, each component of a single tracing step (projection, substepping, tridiagonal linear system solving) takes $O(m)$ time to advance the integration front from $\psi_{j}$ to $\psi_{j+1}$, so tracing $n$ steps is overall $O(mn$). Fitting the surface also takes $O(mn)$ time to build all of the $n$ isoline splines, and $O(n)$ time to evaluate the surface spline at a point (mainly to build $\tilde{\gamma}_{\theta}$). One key advantage of our approach is that the performance is output-sensitive; in other words, it does not strongly depend on the complexity or size of the underlying surface (though in some cases like triangle mesh distance, each implicit distance query is $O(\log N)$ where $N$ is the number of mesh triangles). For example, mesh-based heat methods (Sharp et al., 2019; Herholz and Alexa, 2019) are slightly superlinear in $N$ from solving sparse matrices, but if $mn\ll N$, this is much worse than $O(mn)$. Although the method of Herholz and Alexa (2019) provides fast local solves using only a small output-sensitive subset of vertex values, it still requires a global prefactorization that is superlinear and dependent on $N$ (Herholz et al., 2017). In Table 2, we show a timing breakdown of the examples in Fig. 1, which are representative of the types of implicit functions we used throughout the paper. All timings were recorded on a 2020 MacBook Pro with an M1 processor. From these results, we see that none of the examples took more than 1.5s to complete, and the primary bottleneck is tracing, particularly root finding in each substep (line 8 of Alg. 1). Furthermore, by also looking at the parameters in Table 1 we observe that the terrain point cloud map took roughly twice as long to trace as the brain mesh map, despite the former having 400 times as many vertices as the latter, which confirms the output-sensitive nature of our algorithm. Meshing the log map takes hundreds of milliseconds for every example, though this is because we evaluate the surface at over 10000 points to build the mesh, so this can easily be reduced if a coarser mesh is acceptable. Furthermore, it is apparent that overall performance is proportional to the amount of time required to evaluate the implicit $f$ and its gradient. Overall, these results are sufficient for interactive applications, even on the LiDAR terrain point cloud with over 29 million points, and suggest that speeding up the computation of $f$ and a faster root finding method can further improve performance.
@@ 4.3. Parameter Analysis
The parameter $m$ (number of traced radial curves), $n$ (number of steps taken), and $h$ (step size) are the main quantities used to affect the accuracy of the parameterizations. To demonstrate the effect of these parameters on a simple case with a known analytic solution, we ran our method on a unit sphere (with no substepping and no holonomy smoothing) using a range of values for each of these parameters (where $nh$, the total radius of the map in tangent space, is always $1$) and computed the average $L^{2}$ error of our output from the analytic exponential map, at 2000 samples inside the parameterized disc in tangent space and 500 points along the boundary of the disc (Fig. 7). Overall, we observe that both adding more steps (and correspondingly smaller step sizes) and more radial curves can improve the results, but after enough steps/radial curves, adding more steps/radial curves does not seem to improve the results. In particular, increasing $m$ from 5 to 50 significantly reduces the error, though there is little difference between using 50 or 500 radial curves; as such, we use $m=50$ for most of the results in the paper. Furthermore, we see that the error levels out under refinement of $h$, but at different points when using 5 radial curves (10 steps) and 50/500 radial curves (100 steps). To further investigate the source of the error, we also plotted the error from our traced points along a single radial curve for each step size (since this is on a sphere, there is little difference between each individual radial curve). This error curve follows a very similar trend to the larger $m$ error curves which decrease and eventually level out around $10^{-5}$, which suggests that once there are sufficiently many traced radial curves, the primary source of error is the tracing rather than the spline interpolation. Since the error in the traced points accumulates over more steps, there is a tradeoff between accumulating larger per-step error over fewer steps, and accumulating smaller per-step error over more steps, which causes the error to level out under refinement. Despite the lack of error reduction beyond $10^{-5}$, this is quite small, and as demonstrated in the examples throughout the paper, does not cause practical issues.
@@
Holonomy smoothing strongly influences the output, and is controlled by a regularization parameter $\kappa$. In Fig. 8, we examine the effect of $\kappa$; as it decreases, the effect of the smoothing is diminished, and in the limit as $\kappa\to 0$, the smoothing is completely skipped. We built a local parameterization for a neural implicit representing the Stanford armadillo, around a point near a small indent in the forearm created by isosurface noise. With large $\kappa$ (i.e., strong smoothing), we see that the map has low distortion away from the origin, and although there is some distortion near the pothole itself, this distortion largely disappears as the map expands. Meanwhile, as $\kappa$ decreases, the map’s quality degrades; some regions fold over each other (seen through colormap discontinuities), while other regions significantly stretch away from the origin. As a result, we exclusively use $\kappa=10^{3}$ everywhere else in the paper. However, in large map regions, holonomy smoothing can produce an undesirable shearing effect on the radial curves, where some curves end up becoming nearly parallel, an effect which primarily happens around large surface protrusions (Fig. 9).
Figure 9. In the presence of particularly large surface features, holonomy smoothing produces large rotations and shears the integration front, causing curves to be near-parallel and overlap; the resulting map is also distorted. On this mesh of the log map, the symmetric Dirichlet energy (Smith and Schaefer, 2015) is 6.52, and the least squares conformal map (LSCM) energy (Lévy et al., 2002) is 0.987, both higher than the averages shown in Fig. 11.
@@ 4.4. Comparisons
Compared to prior local parameterization methods, our method can extract significantly more information out of implicit surfaces. Rather than simply sampling the surface for points, and using nearest neighbor information to, e.g., build graphs or Laplacians, we generate the samples we need through tracing, while also being able to evaluate the generated map on other surface points. However, our method does not provide precise control over the size of the resulting parameterization beyond integration parameters, which makes it difficult to directly compare to “log map-based” methods in prior work which require such information as input. That said, our generated maps can still be evaluated on isolated surface samples after the maps have been expanded from a given surface origin, which forms the basis of our comparison methodology against prior work. More specifically, to compare our method to prior work, we generated a local parameterization around a saddle point with 500 traced samples, and rejection sampled the same number of surface points from a 3-dimensional ball around the map origin in $\mathbb{R}^{3}$ such that all the points are contained in our map. Then, we evaluated the log map of our method on those sample points, and also provided those points as input to both versions of discrete exponential maps (DEM) (Schmidt et al., 2006; Schmidt, 2013) and the vector heat method (Sharp et al., 2019). We used our own implementation for the DEM methods and the authors’ released code for the vector heat method on point clouds [^1] https://github.com/nmwsharp/potpourri3d, which uses intrinsic Delaunay triangulations to build more numerically robust Laplacians. For each map, we built a mesh connecting the samples by using the connectivity from the Delaunay triangulation computed in tangent space. As seen in Fig. 10, all other methods produce high distortion away from the origin, and the vector heat method struggles throughout the entire map. It is worth noting that these struggles are exacerbated by the naïve point sampling method, which tends to produce clusters of points that can be prevented by more sophisticated blue noise sampling techniques (Witkin and Heckbert, 1994; De Goes et al., 2012); nevertheless, our method produces low distortion throughout the map despite the sample quality, since we do not depend on them for generating the parameterization.
Figure 11. Across a dataset of surfaces, map positions, and map sizes, almost all of the maps produced by our method have near-minimum symmetric Dirichlet energy, whereas a significant fraction of the maps produced by other methods have higher energy. (top). Our method also produces maps with orders-of-magnitude lower median LSCM energy compared to other methods, though with more variance (bottom). All histograms are plotted on a log scale on the x-axis.
We also used the same point surface point $\to$ log map setup to quantitatively investigate the distortion of our method’s output compared to prior work. More specifically, we used a dataset of over 100 surfaces collected by Myles et al. (2014), and for each surface sampled 10 surface points as map origins, to obtain a set of over 1000 surface regions over which to build local parameterizations. For half of these local surface regions, we sampled 500 surface points within a geodesic radius of 0.1 of the origin, and for the other half we sampled 1000 points within a geodesic radius of 0.2. These points were then evaluated through the log map of each method used in Fig. 10 and triangulated in tangent space to obtain connectivity for the sample points in $\mathbb{R}^{3}$. We used the area-weighted average symmetric Dirichlet energy (Smith and Schaefer, 2015) and least-squares conformal map (LSCM) energy (Lévy et al., 2002) to measure the degree to which the resulting parameterized meshes represent isometries and conformal maps, respectively.
The results of this experiment are provided as histograms in Fig. 11. Across all methods, the symmetric Dirichlet distributions look quite similar, with the majority of maps exhibiting low distortion near the global minimum of 4, and a tail of maps with higher distortion. However, our method (right) produces the largest fraction of low-distortion maps compared to the other tested methods, with over 900 of the maps contained in the lowest histogram bin, and as a result has a much “thinner” tail than the other methods. The DEM methods in particular (left, middle left) have nearly half of their outputs outside of the lowest bin. On the other hand, the LSCM distributions vary significantly across the different methods: DEM methods exhibit a mode near the right side of the distribution near 1, the vector heat method (middle right) has a much narrower distribution but with a lower mode near $10^{-1}$, and our method has a much wider variance but a mode of $10^{-3}$, orders of magnitude lower than the other methods. We can see one such example of a high LSCM energy in Fig. 9, where the map has an LSCM energy of nearly 1. Overall, we see that the combination of geodesic tracing, substepping, and holonomy smoothing produces near-isometric and near-conformal maps in most cases, which quantitatively affirms Fig. 10, but can produce non-conformal results when holonomy smoothing shears regions with large variations in curvature, as seen in Fig. 9.
@@
Figure 12. On a coarse bunny mesh (top), a local parameterization generated by our method (right) avoids the severe artifacts near the boundary of the map (left, centre left) and angular distortion near the origin (centre, centre right) seen by other methods. All methods improve after one level of Loop subdivision (bottom), but the DEM-based methods still possess boundary artifacts (left, centre left). Please zoom in to see map details.
@@
Since we trace points along the surface, another interesting property of our method on triangle meshes in particular is that our result quality is independent of the underlying triangulation. In Fig. 12 we compare our method to both DEM methods, the vector heat method, and another heat-based method (Herholz and Alexa, 2019), for which we used our own implementation, taking care to use intrinsic Delaunay operators from libigl (Jacobson et al., 2018) (though our implementation does not use the accelerated solve, since a public implementation of the specialized prefactorization data structure is unavailable). For this experiment, we computed the 10-ring of an origin vertex (i.e., all vertices at most 10 edges away from the origin) on a coarse Stanford bunny mesh with 3500 vertices, and used these as the evaluation points for all methods. The DEM methods only used these points and their derived connectivity from the mesh, while we used the entire mesh to define the required operators for the heat methods (since we would change the boundary conditions if we only used the local patch). Our method does not build a map from pre-selected discrete points (aside from the map origin), so to make a best-effort comparison we ran our method so that the resulting map both contained the same 10-ring and used a comparable point budget (400 traced samples vs. 453 vertices in the 10-ring — see Table 1 for the parameters we used). The original mesh connectivity was used to connect the 10-ring evaluation points into a continuous patch for all methods. We see that our method produces by far the least distortion in the map, and just like in Fig. 10, the DEM methods exhibit distortion far away from the origin, while the heat-based methods struggle throughout the map, since the triangulation is not of a high enough quality to produce well-conditioned operators. After one level of Loop subdivision (which produces a mesh with roughly 14000 vertices), we ran the same experiment on the 20-ring to cover the same region on the surface (now containing 1720 vertices), and did not change the parameters in our method (though we should clarify that the underlying surface is now the subdivided mesh and not the original mesh). Now we see that our method produces virtually the same result as before, the heat-based methods perform much better with very little distortion, and the DEM methods improve but still exhibit artifacts at the boundary of the map. However, these other methods only achieve this quality with 4 times the number of points that we need to achieve similar quality, and cannot achieve this at all on a coarser mesh.
@@ 4.5.1. Local Texturing
A variety of examples of texturing different representation types is provided in Fig. 1. We show several decals on each surface, where each decal is represented by a meshed version of the spline surface, and can see that none of the resulting images are distorted, while still conforming to the original surface. The terrain point cloud, owl neural implicit, and brain mesh all have challenging geometry with surface bumps, ridges, and divets that make parameterization more difficult. By placing maps in small regions, however, even regions like the owl’s plumage and the heavily folded surface of the brain can be textured.
@@
Figure 13. We obtain high-quality maps on a variety of neural implicit surfaces (rendered as meshes reconstructed with marching cubes), even in the presence of noisy reconstruction artifacts and fine ground truth details. The local maps reveal surface bumps and folds that the marching cubes-reconstructed surface mesh does not capture at its configured resolution.